Hello,
I need help using a system of equations to solve this problem:

Amal reversed the digits in the amount of a check and underpaid a customer by $36. The sum of the digits in the two-digit amount was 8. I need help finding the amount of the check.

Thank you

Answer by scott8148(5891)

let t=(tens digit) and u=(units digit) ... t+u=8 ... t=8-u and u=8-t

10u+t+36=10t+u ... substitute from first line to find t or u; then use t+u=8 to find the other